# A New Perspective on Double-S-Curve Motions of Higher Order and Optimal Motion Planning

### v1

@@ 1 Introduction
Optimal motion planning has been done for many years, and it is still an active research field – particularly from the mathematical – algorithmical and the numerical-computational perspective. Some important applications are robotics (see [1, 2]), aeronautics and astronautics (see [3, 4]), or transport planning (see [5]) – just to mention a few. A rather simple but important task is to find the minimum-time motion for accomplishing a given distance $s$ under certain restrictions: e.g., the velocity is bounded, or the acceleration is bounded, or both are (see [6] for an excellent book about the kinematics of trajectory planning). Common approaches consider the jerk as third derivative as well [7, 8, 9]. Such approaches work in a 7-segment motion (SSM) profile (which is also referred as double-S curve) to find relations between the minimal time, the distance and the boundary values for several cases and under certain boundary conditions.
The aim of this work is to reveal new perspectives on symmetric [^1] For a treatment of symmetric trajectories see [10]. double-S curves of arbitrary order. Following the work of [11] (see also [12]), a general equation for the duration of the motion $T$ is proven. In the case of jerk-controlled motion, this equation is
@@
Here, $v$ denotes the maximum velocity achieved during the motion; $a$ represents the maximum acceleration achieved; and $j$ stands for the corresponding jerk. The generalization to higher order appears quite obvious. Furthermore, it is noteworthy that this Equation (1) holds regardless of how many phases the motion actually consists of [^2] For example, a jerk-controlled motion can consist of four, fove, six or seven phases. Considering the next order (snap), there can occur 15 phases – see [6] where an analysis is done phase by phase..
Moreover, since the equation is true for all symmetric trajectories with vanishing boundary values, it can be used to describe not only the time-minimal motion over the distance $s$. Conversely, if a time horizon and a distance are given, the question arises as to which trajectory – among all possible motions – minimizes, for example, the achieved velocity (or acceleration, or jerk, etc.). Equation (1) and its generalization are also suitable for such questions. In the following, a systematic approach to various optimizations and arbitrary orders is developed using a simple equation.
The paper is organized as follows. Section 2 precisely defines terms such as “symmetric”, which were only used intuitively in the introduction. The class of admissible trajectories is defined and the optimization task under consideration are set up. To familiarize ourselves with the topic, Section 3 deals with the jerk-controlled motion, and Equation (1) is proven for the four possible cases by case distinctions. This is the common approach, done e.g., in [6, 7]. Section 4 presents the general proof for arbitrary orders. Section 5 is devoted to applications with the focus lying on optimal motion planning, especially for jerk- and snap-controlled movements. A brief summary is provided in Section 6
@@ 2 Problem Setup
Let us consider one-dimensional trajectories encoded in $x(t)$ where $x\in\mathbb{R}$ denotes the position as a function of the time $t$ which ranges in the interval $[0,T]$ with $T>0$. In addition, the order parameter is denoted by $N$ with $N\in\mathbb{N}$ and $N>0$. We assume $x(t)$ to be (N1)(N-1)-times continuously differentiable w.r.t. $t$, and the $N$-th derivative of $x(t)$ should be piecewse continuous with countable many finite jumps. Then, $\dot{x}(t)$ is the velocity, and – if existing – $\ddot{x}(t)$ and $\dddot{x}(t)$ stand for the acceleration and the jerk, respectively [^3] The dot indicates derivatives w.r.t. $t$. And in general, the $k$-th derivative w.r.t. $t$ is written as $x^{(k)}(t).$.
@@
A trajectory $x(t)$ is defined to be admissible, if it fullfils the following boundary conditions and constraints:
@@
Cleary, this is nothing else than the classical problem of finding the minimum-time motion for accomplishing a distance $s$ with bounded velocity $(N=1)$, acceleration $(N=2)$, jerk $(N<3)$, $\ldots$
Minimum-Deivative Trajectory (MDT)
@@
Here, both distances $s$ and the time horizon $T$ are given [^4] Clearly, the given time horizon of MDT has to be larger than the one of the minimum-time solution of MTT to ensure the existence of a solution., and the task is to find a trajectory, where a certain derivative (e.g., the velocity if $M=1$) is minimized. Such minimum-velocity motions or minimum acceleration motions (i.e. $M=2$ for $N\geq 3$) are considered because such trajectories protect the device since they avoid wear and tear or large forces for instance.
For both MTT and MDT the key characteristics of the optima are known [9]. Basically, the $N$-th derivative of an optimal trajectory is piecewise constant, and optimal trajectories enjoy a certain kind of symmetry in such a way that – for example – accelerating and decelerating are time-reserved processes. The following definition completes this section by formalizing the above mentioned relevant trajectory type.
@@ 3.1 Acceleration-Controlled Motions
If we consider second-order motions ($N=2$), the acceleration enters the game. Now there are two cases of pre-optimal trajectories depending on the involved parameters $T$, $s$ and $w_{1}$, $w_{2}$: The velocity profile $\dot{x}(t)$ is triangular or trapezoidal (see Figure 1 left and right, respectively). The constant $c$ of Definition 2 (iv) is denoted with $a$ since it has the meaning of an acceleration, i.e. $\ddot{x}(t)\in\left\{-a,0,a\right\}$ for all $t\in[0,T]$ and $a\leq w_{2}$.
@@
This is the ting sister of Equation (1). In Section 5, we turn to the solutions of MTT and MDT based on Equation (6).
@@ 3.2 Jerk-Controlled Motions
By including the jerk constraint $(N=3)$ we now reach the actual case of the SSM. However, there are four cases to be distinguished (see Figure 2). We derive Equation (1) for each of the cases. Let the pre-optimal constant $c$ be $j$, let $v$ be defined as before in Section 3.1, and let $a:=\max\limits_{t\in[0,T]}|\ddot{x}(t)|$ denote the maximal acceleration reached during the motion.
Figure 2: Acceleration profiles $\ddot{x}(t)$ for the four possible cases of a jerk controlled motion.
@@
which is equivalent to $\frac{s}{v}=T-2t_{1}$. We conclude via Equation (7)
@@ 4 Time Horizon of Pre-Optimal Motions
Consider a sequence $(t_{n})_{n=0}$ with $t_{0}>0$ and $t_{n+1}\geq 2t_{n}$ for all $n$. Fix a constant $c_{0}\in\mathbb{R}$ with $c_{0}>0$ and a function $f_{0}:[0,t_{0}]\rightarrow\mathbb{R}$ with $f_{0}(t)=c_{0}$ for all $t\in[0,t_{0}]$. Define for sequences of functions $(\tilde{f}_{n})_{n\geq 0},(f_{n})_{n\geq 0}$ with $\tilde{f}_{n},f_{n}:[0,t_{n}]\rightarrow\mathbb{R}$ and
$\tilde{f}_{n+1}(t)=\begin{cases}f_{n}(t)&for\quad 0\leq t\leq t_{n}\\ -f_{n}(t_{n+1}-t)&for\quad t_{n+1}-t_{n}\leq t\leq t_{n+1}\\ 0&elsewhere\end{cases}$ (33)
$f_{n+1}(t)=\int_{0}^{t}\tilde{f}_{n+1}(\tilde{t})\mathrm{d}\tilde{t}$ (34)
@@
    By induction: The case $n=0$ is clear due to the definition. Consider now $f_{n+1}$ while assuming $f_{n}(t)\geq 0$ for $t\in[0,t_{n}]$. Then $f_{n+1}(t)\geq 0$ for $t\in[0,t_{n}]$ since it is the integral of a positive function. For $t\in\left]t_{n},t_{n+1}-t_{n}\right[,f_{n+1}(t)$ remains constant (and positive) because the integral is 0. For $t\in[t_{n+1}-t_{n},t_{n+1}]$, the absolute value of the negative integral is at most as large as its positive counterpart. Hence, $f_{n+1}(t)\geq 0$ for all $t\in[0,t_{n+1}]$.
@@
    $\int_{0}^{t}f_{n-1}(\tilde{t})\mathrm{d}\tilde{t}$ $=\int_{0}^{t}f_{n-1}(t_{n-1}-\tilde{t})\mathrm{d}\tilde{t}$ (35)
@@
$\int_{0}^{t_{n-1}}f_{n}(t)-\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t$ $=\int_{0}^{\frac{t_{n-1}}{2}}f_{n}(t)-\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t+\int_{\frac{t_{n-1}}{2}}^{t_{n-1}}f_{n}(t)-\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t$ (38)
@@
$-f_{n}(t_{n-1}-t)-\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t$ (39)
[…]
$-\frac{f_{n}(t_{n-1})}{t_{n-1}}\mathrm{d}\tilde{t}$ (40)
$=\int_{0}^{\frac{t_{n-1}}{2}}f_{n}(t)-\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t-\int_{0}^{\frac{t_{n-1}}{2}}f_{n}(\tilde{t})-\frac{f_{n}(t_{n-1})}{t_{n-1}}\tilde{t}\mathrm{d}\tilde{t}$ (41)
$=0$ (42)
@@
$\int_{0}^{t_{n-1}}f_{n}(t)\mathrm{d}t=\int_{0}^{t_{n-1}}\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t=\frac{1}{2}t_{n-1}f_{n}(t_{n-1}).$ (43)
@@
$\int_{0}^{t_{n}}f_{n}(t)\mathrm{d}t$ $=2\int_{0}^{t_{n-1}}f_{n}(t)\mathrm{d}t+(t_{n}-2t_{n-1})f_{n}(t_{n-1})$ (44)
$=t_{n-1}f_{n}(t_{n-1})+(t_{n}-2t_{n-1})f_{n}(t_{n-1})$ (45)
$=(t_{n}-t_{n-1})f_{n}(t_{n-1})$ (46)
@@ 5.1 Velocity-Controlled and Acceleration-Controlled Motions
As mentioned earlier (see beginning of Section 3), the case $N=1$, which is a uniform motion is rather simple since the only relevant equation is
@@
Recall that $v=x_{1}=\max\limits_{t\in[0,T]}|\dot{x}(t)|$ denotes the maximal reached velocity during the movement. Regarding MTT, the obvious optimum is achieved if $v=w_{1}=v_{\max}$ (velocity bound, see Definition 1). Then $T_{\min}=\frac{s}{v_{\max}}$. Conversely, if both distance $s$ and $a$ time horizon $T_{\mathrm{MDT}}$ are given (clearly, $T_{\mathrm{MDT}}\geq T_{\min}$), then the solution of MDT (for $M=1$) is the uniform motion with the minimal velocity given by $v_{\mathrm{MDT}}=\frac{s}{T_{\mathrm{MDT}}}\leq v_{\max}$. Nevertheless, Equation (52) is the very inconspicuous sister of Equation (1).
@@
$\hat{T}=Ta_{\max}^{\frac{1}{2}}s^{-\frac{1}{2}}\qquad\hat{v}=a^{-\frac{1}{2}}s^{-\frac{1}{2}}\qquad\hat{s}=\hat{a}_{\max}=1$ (55)
@@
  • $\hat{v}_{\mathrm{MTT}}=\hat{v}_{\star}$, $\hat{T}_{\mathrm{MTT}}=\hat{T}_{\star}$ is $\hat{v}_{\star}\leq\hat{v}_{\max}$ (triangular case in Section 3.1, see Figure 1 left), or
  • $\hat{v}_{\mathrm{MTT}}=\hat{v}_{\max}$, $\hat{T}_{\mathrm{MTT}}=\frac{1}{\hat{v}}+\hat{v}_{\max}$ otherwise (trapezoidal case, see Figure 1 right).
Figure 3: $\hat{T}$ as a function of $\hat{v}$ (black curve) according to Equation (56) which exhibits a global minimum at $\hat{v}_{\star}=1$ and $\hat{T}_{\star}=2$ (blue cross). The triangular case of MTT occurs if $\hat{v}_{\max}$ lies behind this maximum (green line). Then the minimal time is exactly $\hat{T}_{\star}$. Otherwise (magenta line, trapezoidal case) the optimum of MTT is located at the magenta bullet. The MDT optimum is colored orange. The optimization is colored orange. The optimum is indicated by the orange bullet.
Regarding MDT with a given $\hat{T}_{\mathrm{MDT}}\leq\hat{T}_{\mathrm{MTT}}$ (a horizontal cut-off), the minimal velocity $\hat{v}_{\mathrm{MDT}}$ is given by the smallest solution of the algebraic equation
$\hat{T}_{\mathrm{MDT}}=\frac{1}{\hat{v}_{\mathrm{MDT}}}+.\hat{v}_{\mathrm{MDT}}$ (57)

### v2

# A New Perspective on Double-S Curve Motions of Higher Order and Optimal Motion Planning

@@ 1 Introduction
Optimal motion planning has been done for many years, and it is still an active research field – particularly from the mathematical-algorithmical and the numerical-computational perspective. Some important applications are robotics (see [1, 2]), aeronautics and astronautics (see [3, 4]), or transport planning (see [5]) – just to mention a few. A rather simple but important task is to find the minimum-time motion for accomplishing a given distance $s$ under certain restrictions: e.g., the velocity is bounded, or the acceleration is bounded, or both are (see [6] for an excellent book about the kinematics of trajectory planning). Common approaches consider the jerk as third derivative as well [7, 8, 9]. Such approaches work in a 7-segment motion (SSM) profile (which is also referred as double-S curve, see [10] for early work and [11, 12] as well) to find relations between the minimal time, the distance and the boundary values for several cases and under certain boundary conditions. Fourth order motions are considered in [13, 14].
The aim of this paper is to reveal new perspectives on symmetric [^1] For a treatment of asymmetric trajectories see [15]. double-S curves of arbitrary order. Following the work of [16] (see also [17, 18]), a general equation for the duration of the motion $T$ is proven. In the case of jerk-controlled motion [^2] The term “jerk-controlled” means that the velocity, the acceleration and the jerk are assumed to be bounded., this equation is
@@
Here, $v$ denotes the maximum velocity achieved during the motion; $a$ represents the maximum acceleration achieved; and $j$ stands for the corresponding jerk. The generalization to higher order appears quite obvious. Furthermore, it is noteworthy that this Equation (1) holds regardless of how many phases the motion actually consists of [^3] For example, a jerk-controlled motion can consist of four, five, six or seven phases. Considering the next order (snap), there can occur 15 phases – see [6] where an analysis is done phase by phase..
Moreover, since the equation is true for all symmetric trajectories with vanishing boundary values, it can be used to describe not only the time-minimal motion over the distance $s$. Conversely, if a time horizon and a distance are given, the question arises as to which trajectory – among all possible motions – minimizes, for example, the achieved velocity (or acceleration, or jerk, etc.). Equation (1) and its generalization are also suitable for such questions. In the following, a systematic approach to various optimizations and arbitrary orders is developed using a single equation.
The paper is organized as follows. Section 2 precisely defines terms such as “symmetric”, which were only used intuitively in the introduction. The class of admissible trajectories is defined and the optimization tasks under consideration are set up. To familiarize ourselves with the topic, Section 3 deals with the jerk-controlled motion, and Equation (1) is proven on an elementary level for the four possible cases by case distinctions. This is the common approach, done e.g., in [6, 7]. Section 4 presents the generalization of Equation (1) and the general proof for arbitrary orders. Section 5 is devoted to applications with the focus lying on optimal motion planning, especially for jerk- and snap-controlled movements. In addition, the solution method for finding the minimum-time motions is explained. A brief summary is provided in Section 6.
@@ 2 Problem Setup
Let us consider one-dimensional trajectories encoded in $x(t)$ where $x\in\mathbb{R}$ denotes the position as a function of the time $t$ which ranges in the interval $[0,T]$ with $T>0$. In addition, the order parameter is denoted by $N$ with $N\in\mathbb{N}$ and $N>0$. We assume $x(t)$ to be (N1)(N-1)-times continuously differentiable w.r.t. $t$, and the $N$-th derivative of $x(t)$ should be piecewise continuous with countable many finite jumps. Then, $\dot{x}(t)$ is the velocity, and – if existing – $\ddot{x}(t)$ and $\dddot{x}(t)$ stand for the acceleration and the jerk, respectively [^4] The dot indicates derivatives w.r.t. $t$. And in general, the $k$-th derivative w.r.t. $t$ is written as $x^{(k)}(t).$.
@@
A trajectory $x(t)$ of order $N$ and with time horizon $T$ is defined to be admissible, if it fullfils the following boundary conditions and constraints:
@@
Cleary, this is nothing else than the classical problem of finding the minimum-time motion for accomplishing a distance $s$ with bounded velocity $(N=1)$, acceleration $(N=2)$, jerk $(N=3)$, $\ldots$
Minimum-Derivative Trajectory (MDT)
@@
Here, both distance $s$ and the time horizon $T$ are given [^5] Clearly, the given time horizon of MDT has to be larger than the one of the minimum-time solution of MTT to ensure the existence of a solution., and the task is to find a trajectory, where a certain derivative (e.g., the velocity if $M=1$) is minimized. Such minimum-velocity motions or minimum acceleration motions (i.e. $M=2$ for $N\geq 3$) are considered because such trajectories protect the device since they avoid wear and tear or large forces for instance.
For both MTT and MDT the key characteristics of the optima are known [9]. Basically, the $N$-th derivative of an optimal trajectory is piecewise constant [^6] This is due to Pontryagin’s maximum principle, see [19] for instance., and optimal trajectories enjoy a certain kind of symmetry in such a way that – for example – accelerating and decelerating are time-reserved processes. The following definition completes this section by formalizing the above mentioned relevant trajectory type.
@@
Let $x$ be an admissible trajectory of order $N$ and with time horizon $T$.
@@
Obviously, the constant $c$ in Definition 2 (iv) should be less or equal than the bound $w_{N}$ (see Definition 1 (iv)).
@@ 3.1 Acceleration-Controlled Motions
If we consider second-order motions ($N=2$), the acceleration enters the game. Now there are two cases of pre-optimal trajectories depending on the involved parameters $T$, $s$ and $w_{1}$, $w_{2}$: The velocity profile $\dot{x}(t)$ is triangular or trapezoidal (see Figure 1 left and right, respectively). The constant $c$ of Definition 2 (iv) is denoted with $a$ since it has the meaning of an acceleration, i.e. $\ddot{x}(t)\in\left\{-a,0,a\right\}$ for all $t\in[0,T]$, and $a\leq w_{2}$.
@@
This is the tiny sister of Equation (1). In Section 5, we turn to the solutions of MTT and MDT based on Equation (6).
@@ 3.2 Jerk-Controlled Motions
By including the jerk constraint $(N=3)$ we now reach the actual case of the SSM. However, there are four cases to be distinguished (see Figure 2). We derive Equation (1) for each of the cases. Let the pre-optimal constant $c$ be denoted by $j$, let $v$ be defined as before in Section 3.1, and let $a:=\max\limits_{t\in[0,T]}|\ddot{x}(t)|$ denote the maximal acceleration reached during the motion.
Figure 2: Acceleration profiles $\ddot{x}(t)$ for the four possible cases of a jerk-controlled motion.
@@
which is equivalent to $\frac{s}{v}=T-2t_{1}$. We conclude via Equation (7) that
@@ 4 Time Horizon of Pre-Optimal Motions
Consider a sequence $(t_{n})_{n\geq 0}$ with $t_{0}>0$ and $t_{n+1}\geq 2t_{n}$ for all $n$. Fix a constant $c_{0}\in\mathbb{R}$ with $c_{0}>0$ and a function $f_{0}:[0,t_{0}]\rightarrow\mathbb{R}$ with $f_{0}(t)=c_{0}$ for all $t\in[0,t_{0}]$. Define the sequences of functions $(\tilde{f}_{n})_{n\geq 0},(f_{n})_{n\geq 0}$ with $\tilde{f}_{n},f_{n}:[0,t_{n}]\rightarrow\mathbb{R}$ and
$\tilde{f}_{n+1}(t)=\begin{cases}f_{n}(t)&\mbox{for}\quad 0\leq t\leq t_{n}\\ -f_{n}(t_{n+1}-t)&\mbox{for}\quad t_{n+1}-t_{n}\leq t\leq t_{n+1}\\ 0&\mbox{elsewhere}\end{cases}$ (33)
$f_{n+1}(t)=\int\limits_{0}^{t}\tilde{f}_{n+1}(\tilde{t})\mathrm{d}\tilde{t}$ (34)
@@
    By induction: The case $n=0$ is clear due to the definition. Consider now $f_{n+1}$ while assuming $f_{n}(t)\geq 0$ for $t\in[0,t_{n}]$. Then $f_{n+1}(t)\geq 0$ for $t\in[0,t_{n}]$ since it is the integral of a positive function. For $t\in\left]t_{n},t_{n+1}-t_{n}\right[,f_{n+1}(t)$ remains constant (and positive) because the integrand is 0. For $t\in[t_{n+1}-t_{n},t_{n+1}]$, the absolute value of the negative integral is at most as large as its positive counterpart. Hence, $f_{n+1}(t)\geq 0$ for all $t\in[0,t_{n+1}]$.
@@
    $\int\limits_{0}^{t}f_{n-1}(\tilde{t})\mathrm{d}\tilde{t}$ $=\int\limits_{0}^{t}f_{n-1}(t_{n-1}-\tilde{t})\mathrm{d}\tilde{t}$ (35)
@@
$\int\limits_{0}^{t_{n-1}}f_{n}(t)-\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t$ (38)
$=$ $\int\limits_{0}^{\frac{t_{n-1}}{2}}f_{n}(t)-\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t+\int\limits_{\frac{t_{n-1}}{2}}^{t_{n-1}}f_{n}(t)-\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t$
@@
$\overset{\tilde{t}:=t_{n-1}-t}{=}$ $\int\limits_{0}^{\frac{t_{n-1}}{2}}f_{n}(t)-\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t+\int\limits_{\frac{t_{n-1}}{2}}^{0}f_{n}(t_{n-1})-f_{n}(\tilde{t})-\frac{f_{n}(t_{n-1})}{t_{n-1}}\mathrm{d}\tilde{t}$
$=$ $\int\limits_{0}^{\frac{t_{n-1}}{2}}f_{n}(t)-\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t-\int\limits_{0}^{\frac{t_{n-1}}{2}}f_{n}(\tilde{t})-\frac{f_{n}(t_{n-1})}{t_{n-1}}\tilde{t}\mathrm{d}\tilde{t}$
$=$ $0$
@@
$\int\limits_{0}^{t_{n-1}}f_{n}(t)\mathrm{d}t=\int\limits_{0}^{t_{n-1}}\frac{f_{n}(t_{n-1})}{t_{n-1}}t\mathrm{d}t=\frac{1}{2}t_{n-1}f_{n}(t_{n-1}).$ (39)
@@
$\int\limits_{0}^{t_{n}}f_{n}(t)\mathrm{d}t$ $=$ $2\int\limits_{0}^{t_{n-1}}f_{n}(t)\mathrm{d}t+(t_{n}-2t_{n-1})f_{n}(t_{n-1})$ (40)
$=$ $t_{n-1}f_{n}(t_{n-1})+(t_{n}-2t_{n-1})f_{n}(t_{n-1})$
$=$ $(t_{n}-t_{n-1})f_{n}(t_{n-1})$
@@ 5.1 Velocity-Controlled and Acceleration-Controlled Motions
As mentioned earlier (see beginning of Section 3), the case $N=1$, which is a uniform motion, is rather simple since the only relevant equation is
@@
Recall that $v=x_{1}=\max\limits_{t\in[0,T]}|\dot{x}(t)|$ denotes the maximal reached velocity during the movement. Regarding MTT, the obvious optimum is achieved if $v=w_{1}=v_{\max}$ (velocity bound, see Definition 1). Then $T_{\mathrm{MTT}}=\frac{s}{v_{\max}}$. Conversely, if both distance $s$ and a time horizon $T_{\mathrm{MDT}}$ are given (clearly, $T_{\mathrm{MDT}}\geq T_{\mathrm{MTT}}$), then the solution of MDT (for $M=1$) is the uniform motion with the minimal velocity given by $v_{\mathrm{MDT}}=\frac{s}{T_{\mathrm{MDT}}}\leq v_{\max}$. Nevertheless, Equation (44) is the very inconspicuous sister of Equation (1).
@@
$\hat{T}=Ta_{\max}^{\frac{1}{2}}s^{-\frac{1}{2}}\qquad\hat{v}=va_{\max}^{-\frac{1}{2}}s^{-\frac{1}{2}}\qquad\hat{s}=\hat{a}_{\max}=1$ (47)
@@
  • $\hat{v}_{\mathrm{MTT}}=\hat{v}^{\star}$, $\hat{T}_{\mathrm{MTT}}=\hat{T}^{\star}$ if $\hat{v}^{\star}\leq\hat{v}_{\max}$ (triangular case in Section 3.1, see Figure 1 left), or
  • $\hat{v}_{\mathrm{MTT}}=\hat{v}_{\max}$, $\hat{T}_{\mathrm{MTT}}=\frac{1}{\hat{v}_{\max}}+\hat{v}_{\max}$ otherwise (trapezoidal case, see Figure 1 right).
Figure 3: $\hat{T}$ as a function of $\hat{v}$ (black curve) according to Equation (48) which exhibits a global minimum at $\hat{v}^{\star}=1$ and $\hat{T}^{\star}=2$ (blue cross). The triangular case of MTT occurs if $\hat{v}_{\max}$ lies behind this minimum (green dashed line). Then the minimal time is exactly $\hat{T}^{\star}$. Otherwise (magenta dashed line, trapezoidal case) the optimum of MTT is located at the magenta bullet. The MDT optimization is colored in orange. The optimum is indicated by the orange bullet.
Regarding MDT with a given $\hat{T}_{\mathrm{MDT}}\geq\hat{T}_{\mathrm{MTT}}$ (a horizontal cut-off), the minimal velocity $\hat{v}_{\mathrm{MDT}}$ is given by the smallest solution of the algebraic equation
$\hat{T}_{\mathrm{MDT}}=\frac{1}{\hat{v}_{\mathrm{MDT}}}+\hat{v}_{\mathrm{MDT}}.$ (49)